I really need some help with radicals, could you please check my work?
4/sqrt:3(5x^2) I found this answer 4sqrt:3((5x^2)^2/5x^2)

Answer by stanbon(75887)

4/sqrt[3(5x^2)]
= (4/x)(1/sqrt(15))
Multiply numerator and denominator by sqrt(15) to get:
= (4/x)(sqrt(15)/15)
= (4/15x)sqrt(15)
================
Cheers,
Stan H.